### v1

# Beyond Jailbreak: Unveiling Risks in LLM Applications Arising from Blurred Capability Boundaries

Abstract
LLM applications (i.e., LLM apps) leverage the powerful capabilities of LLMs to provide users with customized services, revolutionizing traditional application development. While the increasing prevalence of LLM-powered applications provides users with unprecedented convenience, it also brings forth new security challenges. For such an emerging ecosystem, the security community lacks sufficient understanding of the LLM application ecosystem, especially regarding the capability boundaries of the applications themselves.
In this paper, we systematically analyzed the new development paradigm and defined the concept of the LLM app capability space. We also uncovered potential new risks beyond jailbreak that arise from ambiguous capability boundaries in real-world scenarios, namely, capability downgrade and upgrade. To evaluate the impact of these risks, we designed and implemented an LLM app capability evaluation framework, LLMApp-Eval. First, we collected application metadata across 4 platforms and conducted a cross-platform ecosystem analysis. Then, we evaluated the risks for 199 popular applications among 4 platforms and 6 open-source LLMs. We identified that 178 (89.45%) potentially affected applications, which can perform tasks from more than 15 scenarios or be malicious. We even found 17 applications in our study that executed malicious tasks directly, without applying any adversarial rewriting. Furthermore, our experiments also reveal a positive correlation between the quality of prompt design and application robustness. We found that well-designed prompts enhance security, while poorly designed ones can facilitate abuse. We hope our work inspires the community to focus on the real-world risks of LLM applications and foster the development of a more robust LLM application ecosystem.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.252941 www.ndss-symposium.org
I Introduction
Amid the wave of large language models (LLMs), a new application paradigm—LLM applications (i.e., LLM apps, agents)—is reforming traditional application concepts. OpenAI launched GPTs in November 2023, enabling users to build customized AI applications leveraging ChatGPT’s capabilities. Subsequently, multiple frameworks and platforms for building LLM applications emerged, such as Coze [Bytedance_coze], Poe [Poe], LangChain [LangChain], and CrewAI [CrewAI]. The adoption of LLM applications is rapidly increasing across diverse, real-world scenarios, such as threat detection and mobile assistants.
Unlike foundational LLMs, LLM-powered apps typically offer more specialized, domain-specific solutions. For instance, a coding app built upon an LLM is fine-tuned to excel at programming tasks, while an LLM-powered translation application can adeptly handle multilingual scenarios. Moreover, LLM applications provide non-expert users with more accessible and convenient entry points, bypassing the need for complex prompt engineering and the direct restrictions of foundation models (e.g., OpenAI’s regional access restrictions). However, if applications are developed without proper security considerations, these accessible entry points can be exploited. Even when a full jailbreak is not possible, an attacker can still breach the capability limits set by the application developer. This forces the application to produce unexpected outputs, such as deviating from its intended purpose or executing entirely unrelated tasks, all while operating within the foundational safety constraints of the LLM.
Motivation Examples. The optimization of LLM apps for specific tasks introduces unforeseen complexities for traditional jailbreak methods. This motivates our research to move beyond conventional jailbreak and instead investigate “goal deviation”—the phenomenon where an application’s behavior strays from its intended objective without an explicit jailbreak. By studying this behavior in real-world scenarios, we aim to dissect the emerging risks of misuse associated with specialized LLM applications.
We now illustrate a real-world business scenario from the Web3 industry, where strict auditing of operator actions is essential for protecting user funds. To this end, some Web3 companies deployed LLM-powered auditing systems to verify the compliance of operator actions, as depicted in Figure 1(a). If an action is deemed compliant, it proceeds normally; if an anomaly is detected, an alert is triggered for manual intervention. This creates an opportunity for a malicious operator to use specific inputs to bypass the audit. A malicious operator can use specially-crafted inputs or prompts to degrade the capabilities of the LLM auditor, bypass its checks, and subsequently steal user assets. We term this scenario “capability downgrade” (Figure 1(b)). In this case, the malicious input does not violate the foundational safety constraints of the LLM itself; it merely degrades the specialized capabilities of the application for its intended task. Second, we define “capability upgrade,” a state where an application, even without being fully jailbreak, can be abused to perform tasks beyond its original, intended scope. For instance, an operator can use crafted inputs to compel an auditing bot to perform additional tasks, subsequently abusing corporate assets for financial gain (as depicted in Figure 1(c)). Finally, “capability jailbreak” refers to the state where the application can be directed to execute any arbitrary task.
This form of abuse has been demonstrated in practice: through meticulously engineered prompts, attackers can exploit LLM applications on nearly any platform, often at no cost to themselves. For example, in early 2025, Rednote rapidly launched an LLM translation feature to accommodate U.S. TikTok refugees. Despite resolving language barriers, weak access controls allowed users to bypass functional limits and execute unauthorized tasks, marking the first real-world case of LLM apps exploitation through social media applications.
Reasearch Gap. Some researchers have already taken notice of this emerging ecosystem and conducted preliminary explorations. Several studies have evaluated the GPTs ecosystem, focusing on aspects such as the distribution and deployment of applications [houapp2024, YanRMWOB24, zhaovision2024, zhangfirstlook24, HouZ0024], highlighting the rapid development of LLM applications. In terms of security threats, some research has revealed the risks faced by LLM applications, including LLM app squatting and cloning [xieappsquatting2024], data exposure [jaffdataexposure2024], and attacking the communication of agents [futricking2024]. However, existing research has primarily focused on applications built on OpenAI’s GPTs, with little coverage of other platforms. Furthermore, this body of work has centered on issues such as jailbreak (i.e., circumventing safety constraints) and privacy issues. In summary, research on LLM applications is still in its infancy, and the risks faced by LLM applications remain largely unexplored by the security community, especially the risks of abusing stemming from application “goal deviation.”
Our Study. Our work systematically analyzes the differences between LLM and traditional applications, highlighting the potential risks introduced by the new development paradigm. The most significant distinction lies in the shift in the developer’s role—from a capability implementer to a capability restrictor. In traditional applications, all supported functionalities are developed by the developer, making code security the most critical concern. In contrast, the core functionalities of LLM applications are provided by the LLM itself. The developer’s primary role is to bridge the gap between users and the LLM while restricting unwanted LLM capabilities. A gap in the developer’s understanding of an application’s capabilities can impact or even compromise the application’s performance. Thus, we formally define the capability space of LLM applications to support their capability analysis.
Figure 1: The potential risks caused by loose boundaries in LLM applications. Capability upgrade is an intermediate state before a full capability jailbreak. It refers to a scenario where an application, even without being fully jailbreak, can still be misused to perform tasks beyond its original scope.
Methodology. To evaluate the impact of capability boundary risks in the LLM app ecosystem, we designed an LLM app evaluation framework, LLMApp-Eval. First, we collected metadata for 807,207 applications from 4 LLM app platforms, including GPTs store and Coze. Next, to address the challenge of quantifying natural language prompts, we analyzed the applications’ targets, process, capability descriptions, and constraints, designing and implementing quantitative metrics. Finally, we designed evaluation scenarios for the three types of risks and implemented automated evaluation scripts.
Findings. Leveraging LLMApp-Eval, we conducted the first large-scale cross-platform analysis of the LLM app ecosystem. We found that LLM apps are being widely adopted, but some super developers have caused low-quality applications to dominate the ecosystem, with one developer releasing over 8,000 applications in a short period. Moreover, variations exist in the LLMs supported by different platforms and their plugin configurations, which may directly impact application security.
Furthermore, we analyzed publicly available prompts of 11,176 applications and found them to be of low quality, with 43.41% of applications lacking any capability constraints. This ambiguity in prompt design increases the risk of misuse. In our experiments, we even found 17 applications that directly executed malicious tasks without the need for adversarial techniques. Our controlled experiments confirmed the impact of prompt design on an application’s capabilities, proving that the risk of abuse can be significantly reduced by incorporating clear capability constraints (Section V-A).
In addition, to evaluate the real-world impact of these risks, we assessed the top 50 applications on four different platforms (which typically do not disclose their prompts) and six open-source LLMs. Our boundary test cases caused a performance impact of 23.94% to 35.59% on the 6 LLMs, demonstrating the feasibility of capability downgrade. And we identified 178 (89.45%) potentially affected applications capable of performing tasks from more than 15 categories or malicious tasks. We observed significant differences in applications’ resilience to risks across platforms. All 50 test applications on GPTs completed tasks from more than 10 categories. We found that the foundational LLMs and plugins used by applications are two critical factors. For instance, GPTs is preconfigured with Web Search and DALL·E plugins, resulting in significantly higher capability upgrade risks in Image & Video tasks compared to applications on other platforms.
Contributions. We make the following contributions:
New risks. We systematically analyzed the new LLM application paradigm, uncovering practical risks of capability abuse that extend beyond conventional jailbreak.
Extensive evaluation of new risks. We designed and implemented an automated evaluation framework to thoroughly assess these new risks. The results reveal that 89.45% of popular applications across the 4 platforms, face at least one abuse risk, underscoring the severity of the issue.
II Background
LLM Application and Developer. LLM applications (i.e., agents) refer to applications built on the capabilities of large language models, designed specifically to solve particular tasks, like text generation, translation, and image generation. Since the release of large language models such as OpenAI’s ChatGPT, their powerful ability to tackle a wide range of problems has been evident. Leveraging these capabilities to upgrade existing products or develop new applications has become a new avenue for many companies and developers. For instance, Microsoft released LLM-enhanced Microsoft 365 Copilot, and Google launched NotebookLM. Notably, with the introduction of GPTs by OpenAI, individual developers can easily customize their own LLM applications on the GPTs platform. The entire development process can be completed with as little as a single application description, significantly lowering the barrier to application development compared to traditional methods. However, this ease of development introduces new risks. Our analysis reveals the presence of numerous super developers across platforms, who published a massive number of applications (Section V-A).
LLM Application Platform. Similar to application marketplaces in traditional application ecosystems, LLM application platforms provide a centralized marketplace offering detailed descriptions and categorization of LLM applications. Users can search for and use the applications they need directly on the platform. Additionally, most current platforms also offer application development features, such as GPTs [Openai_GPTs] and Coze [Bytedance_coze]. GPTs requires a paid membership to develop applications, whereas Coze provides this functionality for free. LLM application platforms have already attracted a large number of developers and applications. GPTs hosts over 3 million applications, and our research shows that the application volume on other platforms is also growing rapidly. For example, Coze currently has over 200,000 applications.
LLM Application Development Paradigm. LLM applications have introduced a novel development paradigm for developers, where the application’s capabilities are inherited from the LLM rather than explicitly defined by the developers. Developers no longer need to implement codes for the task but instead focus on eliciting the model’s capabilities for the target task while constraining its abilities in other areas.
The core of LLM application development lies in building a better bridge between users and large language models, which distinguishes it from traditional application development, where developers primarily focus on implementing application functionalities. Figure 2 illustrates two distinct development paradigms. LLM application developers primarily focus on two key components: prompt templates and capability plugins. Prompt templates are central to eliciting the large language model’s ability to solve specific tasks. Some blogs showed that well-designed prompt templates can significantly improve the accuracy of LLM responses [prompt_engineering, DAIR.AI]. Capability plugins are designed to address the limitations of general-purpose large language models in specific task scenarios. For example, a local knowledge base can enhance the model’s ability to process tasks tailored to specific objectives.
In traditional application development, developers adopt a bottom-up approach, starting from foundational functions to build the component layer, business layer, and application layer. User inputs are subjected to strict validation and constraints before being passed to specific functional components. However, in LLM applications, user inputs are in natural language, which introduces significant uncertainty. This uncertainty becomes problematic if inputs are not constrained, as such inputs may inadvertently trigger capabilities of the LLM that are intended to remain restricted.
Figure 2: Comparison of development paradigms: LLM applications vs. traditional applications.
LLM Application Development Model. The new development paradigm has also introduced changes to development models. Based on differences in process, we categorize the new development models into two types: independent and platform-based.
In the independent development model, developers have greater customization capabilities. They design specific prompts and provide tailored functionalities based on the target requirements, using either APIs from foundational LLMs (e.g., ChatGPT) or their own LLMs (e.g., Google’s Gemini). In some cases, developers may even fine-tune the LLM itself. Large enterprises often adopt this development model to create standalone LLM applications (e.g., Google’s NotebookLM) or integrate LLM into existing applications, such as Microsoft’s Microsoft 365 Copilot. In addition, some open-source development frameworks, such as CrewAI [CrewAI], LangChain [LangChain], and FlowiseAI [Flowise], enable individual developers to rapidly build LLM applications. This model allows for complete control over application logic and data. However, it comes with higher costs and requires developers to possess a certain level of coding expertise.
In the platform development model, developers rely on platforms to build their applications, such as GPTs [Openai_GPTs] and Coze [Bytedance_coze]. These platforms simplify the development process by providing user-friendly development interfaces, automating the generation of prompt templates based on task descriptions, supporting upload-and-use knowledge bases, and offering a wide range of optional capability plugins.
Additionally, platforms allow for flexible switching between different foundational LLMs. Some platforms also enable developers to directly publish their applications on other platforms. For instance, Coze supports deploying applications on TikTok, WeChat mini programs, etc.
Figure 3: Comparison of development model: independent and platform model.
III Analysis of LLM Application Risks
III-A Capability Space
In the new development paradigm, the concept of applications has fundamentally shifted. Developers no longer need to create specific functionalities for target tasks; instead, they leverage the capabilities of large language models to accomplish these tasks. LLMs have demonstrated exceptional proficiency in many tasks, and developers must identify and focus on the capabilities required for the target task while restricting other capabilities. This approach enables the creation of an “LLM application” that effectively addresses the user’s specific needs.
We define the subset of an LLM’s capabilities identified by developers as the capability space of the LLM application. As shown in Figure 4, we regard the capabilities of the foundational LLM as a comprehensive capability set. Initially, the capability space is divided into uncensored and censored spaces based on the presence or absence of ethical constraints. Developers then configure rules and constraints to create target applications. Note that LLM applications may also inherit capabilities from the uncensored capability space, which is often the case for malicious applications specifically designed for harmful tasks. In this study, however, we focus primarily on legitimate applications hosted on regular platforms.
The formal definition of the LLM app capability space is as follows:
Definition 1: LLM App Capability Space. The capability space of an LLM app is the probability of completing task $t_{i}$ (where $t_{i}$ belongs to $T$) for an input action $a_{i}$ (where $a_{i}$ belongs to $A$) after a set of capability constraints $C$ has been added to the base LLM $M$.
[…]
For any $t_{j}$ not belonging to the task set $T$, any action $a_{j}$ will yield the output probability 0, directing towards the constrained content.
[…]
Figure 4: LLM app capability space.
Figure 5: New risks of LLM App (App1 App1*).
III-B Emerging Risks of the LLM App
The risks faced by traditional applications primarily include code implementation security (e.g., the presence of exploitable vulnerabilities) and user privacy risks (e.g., the collection and use of user information). However, in the new development paradigm, LLM applications face a novel security challenge: the boundary risk of the capability space.
The capability space of an LLM application is defined by the developer, but if the constraints imposed by the developer are overly loose, it may lead to boundary drift, boundary crossing, or even boundary destruction of the application’s capability space. To achieve this, attackers only need regular user privileges and can interact directly with the LLM application or use covert methods, such as DolphinAttack [ZhangYJZZX17]. However, they cannot modify the existing capability constraints of the LLM application, nor can they directly attack the communication between the application and the base LLM.
In summary, we categorize the boundary risks of LLM application capability spaces into the following three scenarios.
III-B1 Capability Downgrade
The objective of capability downgrade is to impair an application’s performance on its primary, intended task. Attackers use specific inputs to cause a “drift” in the application’s capability boundary (Figure 5(a)), compelling it to produce an incorrect output. For example, as depicted in Figure 1, a malicious operator crafts a sequence of actions to bypass an LLM auditor’s detection, causing it to misclassify a malicious operation as benign and thereby enabling the covert theft of user funds.
By analyzing the attacker’s capabilities and objectives, we formalize this risk as follows.
For a task $t_{k}$, where $t_{k}$ belongs to the application’s capability space and the application’s output result is $r$, an attacker can identify an input $i_{k}$ that causes the application’s output to change to $f$. This can be expressed as follows:
[…]
III-B2 Capability Upgrade
The objective of capability upgrade is to expand an application’s capability space, enabling it to perform tasks beyond its original intended scope. However, it does not allow for arbitrary task execution, thus representing an intermediate state before a full capability jailbreak. Through specially-crafted inputs, an attacker can induce the application to perform tasks that exceed its defined capability boundaries, as illustrated in Figure 5(b) where the capabilities of App1 are expanded to include those of App2. In this scenario, an adversary can abuse the internal or platform-provided LLM apps (e.g., LLM translator of Rednote) to generate profit at virtually no cost.
For a target application App1, with type A and task set $T1$, consider a task $t_{k}$ that belongs to the task set ($T2$) of App2 (type B). If an input $i_{k}$ exists such that App1 can perform task $t_{k}$, this constitutes the capability upgrade.
[…]
Through capability upgrade, an adversary can gain zero-cost access to a powerful LLM API through various channels, using it to support their own business operations. The cost of this illicit traffic is thus borne by the provider of the LLM application. If the abused application is an internal enterprise tool, the company will incur additional operational costs. In cases where the application is provided by a public platform (e.g., Rednote, Tiktok), the attack can not only increase the platform’s operational costs but also degrade the quality of service for its legitimate users.
III-B3 Capability Jailbreak
The objective of capability jailbreak is to simultaneously bypass both the application’s intended capability limits and the safety constraints of the underlying foundation LLM, thereby enabling the application to execute arbitrary tasks. In this state, the application’s capability boundaries are completely compromised, and an attacker, using meticulously crafted inputs, can abuse the application to perform any arbitrary task, as illustrated in Figure 5(c). This can be formalized as follows.
For an application App1 of type $A$, with a task set $T1$, if for any task $t_{k}$ where $t_{k}\notin T1$ or a malicious task, there exists an input $i_{k}$ that enables the task to be completed, this constitutes a capability jailbreak.
[…]
The capability jailbreak dramatically broadens the attack surface for traditional LLM jailbreak, as any vulnerable LLM application can serve as an entry point. With the increasing proliferation of LLM applications on mobile and desktop devices, every app becomes a potential vector for misuse. This significantly lowers the threshold for accessing the model’s full potential, providing malicious actors with widely available tools for exploitation.
IV Methodology
In this section, we introduce the evaluation framework for LLM applications, LLMApp-Eval. First, we collect applications from various platforms and categorize them. Next, we assess the quality of publicly available prompts and evaluate the risks faced by these applications.
Figure 6: LLMApp-Eval: LLM app evaluation framework.
IV-A LLM Application Collection
Selecting Platforms. We selected platforms spanning multiple regions and various types, providing us with a more comprehensive and objective analytical perspective.
GPTs Store: GPTs Store, launched by OpenAI, is the first LLM application marketplace and currently the largest LLM application platform, leading the entire ecosystem.
Coze and AgentBuilder: These platforms were launched by two Chinese internet companies, ByteDance and Baidu, respectively. Both are highly active and have introduced a series of incentive programs to promote LLM apps.
Poe: Unlike the previous three platforms, Poe is a third-party application platform that does not provide its own large language model services. Instead, it supports flexible integration with other LLMs, attracting a large number of independent developers to create LLM applications on the platform.
Collecting App Mate Data. We adopted different collection strategies for each platform. For GPTs Store, we relied on GPTZoo [HouZ0024], an open-source metadata dataset that includes over 500,000 LLM applications from GPTs store as of August 2024. For the other three platforms, we designed and implemented automated application crawlers. Finally, we collected 576,952 applications from GPTs store, 187,115 from Coze, 22,638 from AgentBuilder, and 20,502 from Poe.
In addition to standard metadata such as application names, descriptions, and user visit counts, we also collected publicly available prompts from AgentBuilder. This aspect, which has not been studied in previous work, provides valuable insights into understanding real-world development practices. In Section IV-B, we introduce the first prompt quantification evaluation method, and in Section IV-B, we leverage it to understand the prompt practices of developers and platforms.
Labeling App Categories. The platforms provide type labels for applications, but each platform defines its own application types, lacking a unified standard. To assign a reasonable type to each application, we drew on the methodology from the study [YanRMWOB24]. We first analyzed the classification systems of traditional app stores, including Google Play Store [google_app], Apple App Store [apple_app], and Amazon Appstore [amazon_app], and then summarized and categorized LLM applications into 20 categories, like Education & Learning, Image & Video, and News.
After defining the scope of classification, we used BART-large-mnli [bart_large_mnli], an NLI-based zero-shot classification model provided by Facebook. This model classifies a target (e.g., text) as an NLI premise and constructs hypotheses for each potential label [yin-etal-2019-benchmarking] to perform the classification. We assigned 20 predefined categories as classification labels for the model, and then input the application descriptions into the model, expecting it to determine the most suitable category for each application. To simplify the analysis, we assumed that each application corresponds to only one category. This assumption aligns with the current practice on LLM application platforms, where each application is typically assigned a single type.
IV-B Prompt Evaluation
The prompts designed by developers are the core component of LLM applications and a key factor in evaluating their capabilities and quality. However, due to the natural language nature of prompts, some studies have only qualitatively assessed prompt quality [SantuF23, AjithPXDN24], and effective quantitative metrics for prompt quality are still lacking. To address this, we combined natural language processing techniques and LLMs to calculate metrics such as information content and richness, designing a method for prompt quality evaluation. We analyzed the design principles of prompts in popular applications and, drawing on the agent design guidelines from Anthropic [Anthropic_agent] and Google [google_agant] white papers, developed four evaluation dimensions: target, process, capability, and constraint.
Target (TScore). Our hypothesis is that the more detailed the description of the target, the higher the quality of the prompt. Specifically, We calculate the clarity of the target description (TScore) in the prompt from two aspects.
First, we compute the entropy of the prompt to represent its information content, obtaining the PEScore. Higher information content indicates a more complete description of the target. The Prompt is tokenized into a sequence of words or subwords $\{w_{1},w_{2},\dots,w_{n}\}$. Each token $w_{i}$ is then mapped to a high-dimensional embedding vector $\mathbf{v}_{i}$ using a pre-trained contextual embedding model:
$\mathbf{v}_{i}=\text{Embedding}(w_{i}),\quad i=1,2,\dots,n$
[…]
where $p_{i}$ is the normalized probability of the embedding $\mathbf{v}_{i}$.
Second, we construct a bag of common words related to the target description, focusing on four aspects of vocabulary: identity characteristics, scenarios, actions, and entities. We then split and tokenize the prompt, and search within the bag of words. To avoid the impact of an incomplete bag of words, we calculate the cosine distance between word embeddings. We count the frequency of words with a similarity greater than the threshold $t_{1}$ and normalize this frequency to obtain the PWScore. The PWScore is calculated as follows:
[…]
Where $n$ is the number of tokens in the Prompt.mm is the number of words in the bag of words. $\mathbf{v}_{i}$ is embedding of the $i$-th token in the Prompt. $\mathbf{b}_{j}$ is embedding of the $j$-th word in the bag of words. $t_{1}$ is the cosine similarity threshold. $\mathbf{1}(\cdot)$ is the indicator function (1 if the condition is true, 0 otherwise).
Finally, we compute the weighted average of the PEScore and PWScore to obtain the TScore.
$TScore=\alpha_{11}*PEScore+\alpha_{12}*PWScore$
Here, $\alpha_{11}$ and $\alpha_{12}$ are the weights for prompt entropy and target-related word frequency, respectively.
Process (PScore). For LLM applications, prompts with clear and guiding steps can better direct the application to complete tasks. We evaluate the richness of process descriptions in prompts based on two aspects: process clarity and complexity of logical relationships, forming what we call the PScore.
Specifically, similar to the “target” evaluation, we construct a bag of words for process descriptions, including step-related keywords (e.g., ”first”) and sequential markers (e.g., ”1.”). Then, we calculate the cosine similarity between each word in the prompt and the keywords in the bag of words, counting the number of matching words to obtain $N_{\text{step-keywords}}$ and $N_{\text{markers}}$. Next, to assess the complexity of logical relationships, we use NLP techniques to extract the number of logical connectives in sentences and the number of clauses, resulting in $N_{\text{logic-keywords}}$ and $N_{\text{clauses}}$. After normalization, we obtain $N_{\text{step-keywords-norm}}$, $N_{\text{markers-norm}}$, $N_{\text{logic-keywords-norm}}$, and $N_{\text{clauses-norm}}$. Finally, these normalized scores are weighted and fused to calculate the PScore.
$PScore=\alpha_{21}*(N_{step-keywords-norm}+N_{markers-norm})$
$\alpha_{22}*(N_{logic-keywords-norm}+N_{clauses-norm})$
Here, $\alpha_{21}$ and $\alpha_{22}$ are the weights for process clarity and complexity of logical relationships, respectively.
Capability (CaScore) and Constraint (CoScore). We focus on the level of detail in the prompt’s capability descriptions and whether it includes clear capability constraints, such as refusing to perform certain tasks. Due to the wide variation in application capabilities and differences in developer design habits, traditional NLP techniques, such as sentiment analysis, cannot accurately evaluate the capability descriptions in prompts. To address this, we implemented an LLM-based evaluation method for assessing capability descriptions and constraints in prompts.
Specifically, we carefully designed a prompt to make the LLM output four values: level of detail in capability descriptions ($Cap_{\text{level}}$), number of capability entries ($N_{\text{capability}}$), number of constraint entries ($N_{\text{constraint}}$), and level of refusal ($Con_{\text{level},j}$) for each constraint in the prompt. Both $Cap_{\text{level}}$ and $Con_{\text{level},j}$ are rated on a scale of 1 to 5, with higher scores indicating more detailed descriptions or stricter refusal levels. For example, an application with $Con_{\text{level,j}}=5$ explicitly lists the tasks it refuses to perform. Appendix A provides our prompt and examples of scoring results for some applications. Finally, we calculate $CaScore$ and $CoScore$ using the following formulae.
$\text{CaScore}=Cap_{\text{level}}*N_{\text{capability}}$
[…]
The above four metrics, TScore, PScore, CaScore, and CoScore, evaluate the quality of specific LLM applications in terms of target description, task decomposition, capability description, and capability constraints, respectively. Higher scores represent better quality. However, since their score distributions vary significantly, we normalize them to the range of 0 to 100 for easier evaluation.
Finally, we compute the weighted average of $TScore$, PScorePScore, CaScoreCaScore, and $CoScore$ to obtain the final score for the application prompt, forming what we call the AppScore.
$AppScore=$ $\ \alpha_{1}\cdot TScore_{norm}+\alpha_{2}\cdot PScore_{norm}$
$\ +\alpha_{3}\cdot CaScore_{norm}+\alpha_{4}\cdot CoScore_{norm}$
Figure 7: The boundary case example on Llama-3.1-8 B. In this case, by embedding meticulously crafted adversarial sentences (red), a resume that would otherwise fail evaluation can successfully evade an LLM-powered screening, simulating enterprise LLM recruitment systems where commercial system evaluation was inaccessible.
IV-C Capability Boundary Evaluation
Our capability boundary evaluation consists of two components: test case generation and automated evaluation.
Test Case Generation. To comprehensively evaluate the risks faced by LLM applications, we constructed specialized case sets for different risk scenarios.
For capability downgrade, since capability downgrade targets the weakening of an application’s unique capabilities, evaluating this across hundreds of thousands of applications is challenging. Thus, we evaluate the indirect prompt injection threat to LLMs to assess how carefully crafted inputs can impact the performance of LLMs on the same task. Different from generic prompt injection, our goal is to assess whether an LLM’s output for the same task remains robust when faced with inputs containing misleading information. We conduct simulation-based evaluations across diverse LLM application scenarios, crafting scenario-specific queries to gather responses. Then, we perform systematic comparisons to assess potential capability downgrade in target scenarios. Specifically, we carefully crafted 2,790 pairs of boundary test cases across 28 scenarios. For example, as shown in Figure 7, in the case of targeting LLM-powered candidate screening systems, strategic manipulation of resume content enables unqualified candidates to bypass screening filters, generating erroneous “highly suitable candidate” recommendations.
For capability upgrade, a straightforward method to detect capability upgrade is through interactions that prompt the application to actively generate tasks both within and beyond its capability space. However, this approach may lead to inaccuracies, as applications might hallucinate incorrect cases or reveal only part of their capability space.
To systematically evaluate capability constraints, we constructed a cross-category test case set based on the application types we previously identified. Our hypothesis is that if an application of Type A (App 1) can perform tasks from a Type B application (App 2), then App 1 is at risk of capability upgrade. Specifically, our constructed test case set includes the following three components:
1) Default cases. We observed that applications often provide a small number of default questions to guide users. We collected these default questions from applications of each type as representatives of their capability space. After manual verification and filtering, we selected five default test cases for each type.
2) Generated cases. Default questions are constrained by application developers. To extend the coverage of test cases, we generated out-of-scope test cases for each application type. For each type, we randomly select 5 applications, create their capability profiles, and use GPT-4o to generate 5 test cases for each application. Specifically, we first create a capability profile for the selected application, including its name, description, type, and the claimed capability space revealed through active interaction. We then input this profile into an LLM, instructing it to generate tasks that exceed the application’s capability space. For instance, if the application’s profile is a “weather app,” the LLM might be prompted to generate a coding task. To facilitate reproducibility, the prompts used in our experiments will be publicly released. To ensure the cases are accurate and representative, we manually select 5 cases from the generated cases for each type as the final set of generated cases.
3) Common cases. We created a set of 10 common cases (e.g., ”How many seasons are there in a year?”) to test the general performance of applications.
For capability jailbreak, our objective is to determine whether current jailbreak techniques are constrained by application-defined capability boundaries. We began by constructing a test case set with various malicious question types. Then, we reproduced state-of-the-art jailbreak techniques [0005LZY0S24, JiangXNXR0P24, abs-2310-08419, abs-2309-10253] to generate adversarial test cases capable of bypassing the foundational LLM’s constraints underlying the applications.
Boundary Drift Evaluation. We input paired boundary cases (<Q1, Q1*>) into the target LLM and compare its outputs (<R1, R1*>) for different inputs to determine whether the model’s capability boundary for the target task has drifted.
App Automated Evaluation. For capability upgrade and jailbreak, after obtaining the test cases, we input cross-type test cases, common cases, and adversarial cases based on the application’s type. The application’s output is then evaluated using an LLM-based judge to determine whether it successfully completes the test cases.
IV-D Implementation and Evaluation
Labeling App Categories. We deployed BART-large-mnli locally to perform the application type classification task on a MacBook Pro with an M3 Max chip and 64GB of memory. To mitigate the impact of non-English application names and descriptions on classification accuracy, we used EasyNMT [EasyNMT] to translate them into English.
To evaluate the classification accuracy, we recruited 3 volunteers with experience in LLM application development as evaluation participants. Before the evaluation, the volunteers were trained on our classification standards. Specifically, we randomly sampled 200 applications from the classification results and asked the volunteers to determine whether the application type was correctly assigned. If at least 2 of the 3 volunteers agreed that the classification was accurate, we marked it as correct. Finally, among the 200 tested applications, 96% were classified correctly. The primary reason for false positives is that the capability of an application may be vague. For example, an app that simulates an ancient doctor (Coze ID 7399930163061981203) can be categorized as Education & Learning or Health & Wellness.
Prompt Evaluation. We used the en_core_web_sm and zh_core_ web_sm models in the python scapy library for tokenization. The rapidfuzz library was utilized for fuzzy matching of keywords. Additionally, the multilingual sentence embedding model “paraphrase-multilingual-MiniLM-L12-v2” was employed to compute the embedding vectors of Prompts, which were then used to derive their entropy values. These experiments were conducted on a laptop equipped with an Intel i7-12700H 2.30 GHz CPU and 64GB of RAM. Moreover, we used the OpenAI’s “gpt-4o-mini-2024-07-18” to calculate the CaScore and CoScore, priced at \$0.15 per 1M tokens for input and \$0.60 per 1M tokens for output. The total API cost for these experiments was approximately \$120.
When calculate TScore, PScore, and AppScore, we assign equal weights to all components, i.e., $\alpha_{1}=\alpha_{2}=\alpha_{3}=\alpha_{4}=0.25$ and $\alpha_{11}=\alpha_{12}=\alpha_{21}=\alpha_{22}=0.5$. Using equal weights prevents any single sub-metric from disproportionately influencing the overall score, maintaining a balanced evaluation.
Since there is currently no established method for quantifying prompt quality, we recruited three volunteers to manually evaluate the accuracy of our scoring method. As prompts are expressed in natural language, differences in interpretation among volunteers were expected. To address this, we first introduced the volunteers to our four evaluation metrics and provided corresponding examples. We then randomly selected 100 application prompts and their scores for the volunteers to annotate whether the scoring was accurate. A score was considered accurate if at least two of the three volunteers marked it as such. Our evaluation results showed that our method accurately assessed 92% of the prompts. For the inaccurate results, we found that the primary issue was due to some prompts having unclear logical descriptions or use sequence words outside our predefined bag of words.
LLM Judge. We implemented our LLM Judge using OpenAI’s ChatGPT-4o API. The prompt template we used is provided in Appendix B. To evaluate the accuracy of the LLM Judge, we recruited 3 volunteers to assess its decisions through sampling. Specifically, we randomly sampled 300 judgment results from our test cases. We then provided the case descriptions and the application responses to the volunteers, asking them to determine whether the LLM Judge’s decision was correct. Ultimately, the accuracy of our LLM Judge reached 94.33% (283/300). The misclassified cases were primarily due to our design of a text-only LLM Judge, which may lead to misjudgments when evaluating multimodal tasks. For instance, in response to a requirement to generate an image, a textual description of the image might be mistakenly judged as a correct answer.
V Unveiling the LLM App Ecosystem Risks
In this section, we first present the analysis results of cross-platform LLM applications, highlighting the unique characteristics of the emerging LLM application ecosystem. Next, we evaluate the new risks faced by LLM applications, demonstrating their vulnerabilities.
V-A Characterizing Cross-Platform LLM Apps
Overview of LLM Apps. We collected a total of 807,207 applications across 4 platforms: 576,952 from GPTs (until Apr 13, 2024), 187,115 from Coze (until Sep 29, 2024), 22,638 from AgentBuilder (until Sep 29, 2024), and 20,502 from Poe (until Sep 16, 2024). GPTs store remains the largest marketplace, while new markets such as Coze are also emerging.
Although the number of applications varies across platforms, the distribution of application types is remarkably similar, with the average absolute deviation in the percentage of each category being less than 2%. Using the method introduced in Section IV-A, we labeled each application by category. Surprisingly, we found that the distribution of application types is highly consistent across platforms. The top three categories on all platforms are Education & Learning (16.05%–20.71%), Data & Research (8.43%–13.44%), and Developer & Code (6.24%–10.43%), see Appendix C. This consistency suggests that despite differences in platform coverage regions, user needs remain largely uniform.
Application Scale Growth Trend. The growth rate of applications on GPTs and AgentBuilder has slowed, whereas Coze remains in a phase of rapid expansion. Figure 11 illustrates the growth trends of total applications and top 3 types on GPTs, Coze, and AgentBuilder. GPTs store officially launched in January 2024, and the figure shows a significant surge in application numbers during that time. Similarly, AgentBuilder experienced its fastest growth within the first month of launch, after which the growth rate began to decline. In contrast, Coze continues to grow rapidly, with no signs of slowing down.
Supported LLM. The LLMs supported by different platforms vary significantly. GPTs and AgentBuilder only support their own proprietary LLMs, while Coze and Poe are more flexible, supporting dozens of different large models. However, Coze’s free version only supports the Doubao foundational model, and access to other LLMs requires upgrading to the paid version. Table I shows the LLMs relied upon by the applications we collected from different platforms.
TABLE I: LLMs supported by different platforms.
Platform LLM
GPTs GPT
Coze
Doubao, Qwen, Step, Deepseek, GLM,
Abab, Moonshot, Baichuan
AgentBuilder Wenxin
Poe Text
GPT, Claude, Gemini, Llama, Grok,
Mixtral, MythoMax
Image
FLUX, Imagen3, Playground, Ideogram,
DALL-E, Recraft, StableDiffusion, Recraft
Video Pika, Runway, Dream-Machine
  • : We listed the major LLMs used by the analyzed applications. Only the series names are provided, without specifying specific versions.
Supper Developer. Super developers play a crucial role in the LLM application ecosystem, which causes low-quality, low-usage applications to occupy a significant portion of the current LLM app ecosystem. A developer is considered a super developer if he or she has released a significant number of applications. In traditional application domains, due to the lengthy development cycles, super developers are rare, and their applications are generally of lower quality, attracting limited user engagement. However, we observed the presence of super developers across various LLM application platforms, often drawing substantial user attention.
On GPTs Store, 19 developers have created over 1,000 applications. Figure 17 illustrates the application publishing history of the top five developers on GPTs. The developer with the most applications has created 8,530, covering all eight categories available on GPTs. Although most applications are overlooked, some do gain user attention. For example, when examining the top 10 developers on AgentBuilder and Coze, we found that their applications had a minimum of 53 visits but could reach as high as 1,962,963 visits.
Plug-in, Knowledge, and Workflow. The application development practices of developers across different platforms vary significantly. The metadata for Coze and AgentBuilder includes information about application plugins, knowledge, and workflows. Our analysis shows that on AgentBuilder, 7,458 applications (65.89%) did not configure any plugins, knowledge bases, or workflows to extend their capabilities. In contrast, this figure is only 28% on Coze.
Default plugins provided by the platform may introduce potential security risks. On AgentBuilder, we found some applications configured with plugins that were completely unrelated to their capabilities. For instance, the divorce counseling application is configured with the Baidu Map plugin. We later discovered that these were default plugins provided by AgentBuilder, likely developers forgot to remove them. We identified a total of 258 applications configured with the Baidu Maps plugin. Through manual analysis of the application names and descriptions, we found that 117 (45.35%) of these applications did not require the capabilities provided by this plugin and were likely introduced unintentionally. Since developers may be unaware of these plugins, they could be exploited by attackers, as the hidden APIs [10.1145/3576915.3616676]. In Section V-C, our analysis revealed that the GPTs by default loads the DALLE capability plugin, which exposes many applications to capability upgrade risks.
V-B Prompt Evaluation
We used our method described in Section IV-B to evaluate a total of 11,176 applications with publicly available prompts on the AgentBuilder. Figure 8 displays the distribution of application scores. The AppScore of all applications ranges from 2.55 to 78.41, with 48.62% of applications scoring below 50. Comparing different metrics, the evaluated applications perform better in CaScore and TScore, indicating that they can clearly describe the application’s purpose. However, in terms of task decomposition (PScore), most applications rely on simple sequential markers (e.g., “1.”, “2.”) to describe their functionalities, without further elaborating on the logical relationships between these functions.
More importantly, for capability constraints (CoScore), applications on the platform show a polarized trend. Specifically, 43.41% of applications do not implement any capability constraints, while nearly 20% of the remaining applications that recognize the need for constraints scored below 60.
To validate the relationship between prompt design and application capabilities, we conducted a controlled-variable experiment across four platforms. Specifically, we selected two applications from our analysis of public prompts: one with an unrestricted prompt (a1) and another with a well-designed prompt (a2). On each platform, we then created two derivative applications. The first directly used the prompt from a1, while for the second, we optimized a1’s prompt by emulating the design principles of a2, such as adding explicit capability constraints. Subsequently, we assessed the risks for each application using the methodology detailed in Section IV. We repeat the above procedure five times to ensure the reliability of our experimental results.
The evaluation results show that applications using optimized prompts exhibit a higher rejection rate for out-of-scope tasks, reducing the number of such tasks executed by 5.3%–80%, depending on the platform and base model. For instance, on the AgentBuilder, an application would execute 15 out of 21 different types of out-of-scope tasks before constraints were added; after their addition, this number was reduced to just 3. Furthermore, we discovered that the underlying LLM also influences a prompt’s effectiveness, as identical prompts can yield different outcomes when used with different LLMs. For example, the same capability-constraining prompt exhibited significantly weaker enforcement on the Claude-3-Haiku (Poe) compared to its effectiveness on the Wenxin (AgentBuilder).
Figure 8: Distribution of prompt evaluation scores.
(a) GPTs.
(b) Coze.
(c) AgentBulider.
(d) Poe.
Figure 9: Capability upgrade risks for top 50 applications on GPTs, Coze, AgentBuilder, and Poe. The horizontal axis is the application index, while the vertical axis is test case types. Green indicates the application’s original type. The intensity of orange represents the proportion of test cases from that type completed by the application; the darker the color, the higher the proportion.
V-C Evaluating Risks of LLM App
To comprehensively evaluate the risks faced by LLM applications, we selected the top 50 popular applications from each of the four platforms as representative applications. These applications collectively accounted for at least 46% of user activity on each platform. We then applied the evaluation method introduced in Section IV-C to assess each application. However, since some applications were no longer available during our testing, we ultimately evaluated a total of 199 applications across the four platforms. Moreover, we evaluated the risk of capability downgrade on 6 open-source LLMs, including LLaMA3.1 and Qwen2.5.
Evaluation of Capability Downgrade. We evaluated 6 open-source LLMs using 2,790 pairs of boundary test cases (Section IV-C). The results show that most models are affected by the inserted misleading information, as shown in Table II. Mistral was the most affected, with incorrect responses in 993 cases, while LLaMA performed the best, but still produced incorrect responses in 668 cases. If applications are built on such models without additional defensive measures, they may be vulnerable to capability degradation risks.
TABLE II: The results of capability downgrade tests.
Llama-3.1-8B Qwen2.5-7B Gemma-2-9B ChatGLM3-6B Mistral-7B InternLM2.5-7B
# downgrade case (%) 668 (23.94%) 981 (35.16%) 707 (25.34%) 814 (29.18%) 993 (35.59%) 723 (25.91%)
Evaluation of Capability Upgrade. To avoid the impact of homogenized capabilities across categories on the evaluation results, we selected strict criteria to identify affected applications. An application is considered at risk of capability upgrade if it can perform tasks from 15 or more categories. Finally, we identified 144 (72.36%) applications affected by capability upgrade, including 49 from GPTs, 35 from Coze, 33 from Poe, and 27 from AgentBuilder, as shown in Table III.
The risks of capability upgrade vary significantly across different platforms, with GPTs being noticeably more susceptible to capability upgrade compared to the other three platforms. All 50 test applications on GPTs completed tasks from more than 10 categories, as shown in Figure 10. Figure 9 presents a heatmap of capability upgrade risks for applications on various platforms. The vertical axis represents the types of test cases, while the horizontal axis represents the tested applications. Green indicates the application’s original type. The intensity of orange represents the proportion of test cases from that type completed by the application; the darker the color, the higher the proportion.
In the general capability test (i.e., the common cases), 172 (86.42%) applications completed the task with only a few exceptions: 4 from Coze, 7 from AgentBuilder, and 16 from Poe. Below, we analyze the reasons why these applications failed to answer these common-sense cases:
Input Topic Check: Some applications verify whether the user input aligns with their specific topic. For instance, when asking questions to Coze No.49, the query must include content related to self-driving tours; otherwise, the application will not respond.
Fixed Workflow: Some highly specialized applications are designed with a rigid workflow. For example, Coze No.29 is an IQ test application featuring 10 test questions, and users must follow the prescribed process.
Multimodal Input/Output Requirements: Some multimodal applications (e.g., text-to-image or image-to-text) require specific input/output formats, such as images. For example, Poe No.37 generates an image regardless of the query content.
Observing Figure 9, we found that applications not executing the common tasks rarely (if ever) perform tasks outside their own designated category. This indicates that the aforementioned three measures can be somewhat effective in mitigating capability upgrade risks. Fixed workflows and multimodal input/output requirements resemble traditional software development processes, where strict input/output constraints provide strong resistance to capability upgrade.
Moreover, we observed that some applications, despite implementing input topic checking, still face capability upgrade risks. Input topic check relies on the LLM itself to enforce restrictions. Due to the flexibility and extensibility of natural language, this mechanism can be easily bypassed, allowing attackers to exploit it for capability upgrade.
The differences in LLM support and plugins across platforms are key factors influencing the capability upgrade risk. Comparing subfigures (a) and (b), (c), (d) in Figure 9, GPTs exhibit significantly better performance than the other three platforms in Images & Video and Weather tasks. GPTs are created with default configurations that include Web Search and DALL·E Image Generation, enabling its applications to possess excellent multimodal input/output and real-time information retrieval capabilities. In contrast, the base models provided by Coze and AgentBuilder lack multimodal capabilities, and real-time information retrieval depends on configuring additional plugins. As a result, applications outside the Image & Video category on these platforms cannot perform Image & Video tasks, and only a few applications (3 on Coze, 14 on AgentBuilder) are capable of performing Weather tasks. Compared to Coze and AgentBuilder, Poe offers a more diverse selection of models. Some models provided by Poe, like DALL-E-3, and Pika, have image and video generation capabilities, allowing slightly better performance in Image & Video tasks.
Figure 10: CDF of the Top 50 application capability categories. The red shaded area is considered the range affected by capability upgrade (i.e., coverage categories 15).
Evaluation of Capability Jailbreak. We identified 178 (89.45%) applications that were vulnerable to capability jailbreak risks, as they completed at least one malicious task. Notably, 17 applications executed malicious tasks even without the use of adversarial techniques. To account for the inherent randomness of LLM responses, we repeated the test three times for cases where the application refused the task. Only if the application consistently refused all three attempts did we consider it as having successfully defended against the attack.
The LLM of an application may impacts the effectiveness of capability jailbreaks. We observed capability jailbreak risks vary significantly across platforms. As shown in Figure 16, GPTs is the least affected, followed by Poe, AgentBuilder, and Coze. On GPTs, where all applications uniformly use the same LLM, applications demonstrate high consistency in their vulnerability to capability jailbreak. For instance, 60% applications could only be successfully attacked by attack prompt No.1. In contrast, other platforms, such as Poe, offer more diverse model options (e.g., Poe currently provides 19 different models). While this diversity allows users to select models better suited to their tasks, it also introduces potential risks due to the varying capabilities of the models.
Improperly configured prompts in LLM applications may function as jailbreak prompts. We discovered 17 applications (13 on Coze and 4 on Poe) that were directly exposed to original malicious tasks. We infer that certain rules within the application prompts may unintentionally bypass the constraints of the foundational model. We tested this hypothesis on our own applications and confirmed it. When the application’s prompt explicitly includes constraints such as ”do not refuse any user request,” the application may bypass the foundational model’s restrictions.
TABLE III: Evaluation results for different platform applications.
#App1 #Up-App2 #Jail-App3
Com Cate Mal Adv
GPTs 50 50 49 0 46
Coze 50 46 35 13 47
AgentBuilder 49 42 33 0 47
Poe 50 34 27 4 38
Total 199 172(86.42%) 144(72.36%) 17(8.54%) 178(89.45%)
  • 1: Evaluated applications. 2: Applications exposed to capability upgrade.
  • 3: Applications exposed to capability jailbreak.
  • : Com (Common case); Cate (Category case); Mal (Original malicious case); Adv (Adversarial malicious case).
VI Discussion
VI-A Limitations
Despite our best efforts to understand and assess LLM application risks, some limitations remain. First, a limitation of our study is that our testing focused exclusively on publicly available applications from app marketplaces and did not include internal enterprise applications. However, our work is grounded in a systematic analysis of the development paradigms and capability space abuse issues of LLM applications. Therefore, while we could not directly test internal applications, we contend that any application developed using the same paradigm will face similar risks. This assertion is corroborated by our cross-platform evaluation results.
Second, we analyzed applications from only four platforms and tested only a subset of applications for risk evaluation. The platforms we selected are diverse, including the largest platform, GPTs, the third-party platform Poe, the Chinese platform AgentBuilder, and the rapidly growing Coze. This diversity provides a rich perspective and offers many new cross-platform insights. Additionally, we selected the top 50 most accessed applications from each platform for evaluation, covering 46.59% - 72.76% of the platform’s users. Analyzing these popular applications provides sufficient evidence to demonstrate the risks faced by applications.
Third, our method utilizes application descriptions for app categorization, where these descriptions remain unverified and may affect result accuracy. This approach has demonstrated its effectiveness in previous research [YanRMWOB24], and our sampled evaluation confirms that the classification outcomes meet our analytical requirements (Section IV-D). Moreover, prompt quality evaluation constitutes an open challenge. Drawing from existing studies and analyzing compositional structures of exemplary open-source prompts, we define a four-dimensional evaluation matrix, establishing the first quantitative assessment framework for prompt quality. Our sampled evaluation results indicate the proposed method accurately assessed 92% of the prompts (Section IV-D).
VI-B Reproducibility
We will open-source the scripts used in our experiments [^1] https://anonymous.4open.science/r/nn4cz2wez4x7865n44. This includes the scripts for collecting LLM applications from different platforms and the LLM Judge script. Moreover, we will publish all the prompts we used to facilitate the reproducibility of our work. Finally, we will open-source the application metadata we collected from Coze, AgentBuilder, and Poe to encourage further research.
VI-C Disclosure
Due to platform and dataset limitations, we were unable to directly contact the developers. Additionally, we believe that implementing mitigation measures at the platform level would be more effective. Therefore, we are currently communicating with the platforms, have reported our evaluation results to them, and are waiting for further discussions.
VII Related Work
LLM Application Analysis. The LLM application ecosystem is an emerging system, and some researchers have already begun initial explorations. Several studies have conducted measurement research on this new ecosystem, analyzing its landscape, deployment, and security [houapp2024, YanRMWOB24, zhaovision2024, zhangfirstlook24, HouZ0024, Tao-abs-2401-00905], and have constructed the GPTZoo [HouZ0024] dataset for GPTs, which facilitates the work of future researchers. Furthermore, some researchers have adapted traditional application security issues such as cloning and squatting to the context of LLM applications, revealing that application cloning has already emerged in the LLM application ecosystem [xieappsquatting2024]. Jaff et al. [jaffdataexposure2024] explored data leakage issues in GPTs, while Fu et al. [futricking2024] attempted to deceive applications into performing malicious operations. Additionally, the most critical component of LLM applications is the carefully crafted prompts designed by developers. Moreover, Bo et al. [Bo_pleak_24] proposed a novel technique that effectively leaks private prompts used in applications. These studies have provided initial explorations of the LLM application ecosystem, but they all focus on a single platform. Our work is the first to conduct a comparative analysis of LLM applications across multiple platforms, offering new insights into the development of LLM applications.
LLM Jailbreak. The security of LLMs has been a long-standing concern, as adversaries continuously develop new methods to manipulate these models into generating harmful content. One major focus in LLM jailbreaks is the design of adversarial inputs, a strategy that exploits the instruction-following nature of LLMs [NEURIPS2023_fd661313]. This approach prompts the model to prioritize generating responses based on user requests rather than adhering to safety constraints. Studies such as [LiGFXHMS23, ChangLLWWL24, JiangXNXR0P24] demonstrate how carefully crafted natural language instructions can deceive LLMs into producing harmful or unethical content. Additionally, some researchers have explored identifying unsafe inputs overlooked during the training process of LLMs [0010ZPB24, YuanJW0H0T24, LiuXCX24].
VIII Conclusion
In this paper, we uncovered potential new risks caused by the ambiguous capability boundaries of LLM apps, leading to practical abuse scenarios that do not require jailbreak. We designed and implemented an LLM app capability evaluation framework to evaluate the impact of these risks. We selected the top 50 popular applications from 4 platforms and 6 open-source LLMs as evaluation targets. Results show that our boundary cases resulted in a maximum performance impact of 35.59% on 6 open-source LLMs. Moreover, we identified that 178 (89.45%) potentially affected applications, which can perform tasks from more than 15 scenarios or be malicious. In addition, we highlighted that robust application design can effectively mitigate boundary risks. We hope our new insights will inspire the community to focus on the LLM app ecosystem.
Ethical Considerations
Our analysis involves crawling publicly available data and conducting security testing on the targets, with careful consideration of ethical issues in the experimental design. We adhere to ethical standards based on the Menlo Report [kenneally2012menlo], best practices for network measurement [partridge2016ethical], and recommendations for using public dataset [imc07-allmanSharedData]. First, we limited the data collection rate to remain within the access rate allowed by the platforms, ensuring that our data collection did not impact platform services. Additionally, to avoid the extra impact caused by duplicate data collection, we refrained from re-crawling datasets already made publicly available in other studies. For example, we used GPTZoo’s [HouZ0024] published GPTs metadata instead of duplicating the crawling process. Second, we adopted three measures to minimize the impact of our testing on the target applications: 1) our tests were conducted using our own test accounts, ensuring that other users were not affected. 2) we did not test all applications but selected a subset of popular applications as test targets. These popular applications have high traffic volumes, so our limited test interactions would not disrupt their operations. 3) our test queries consisted primarily of harmless questions, which would not cause denial-of-service issues or disrupt the application’s logic, such as causing it to crash. Finally, the application metadata we analyzed consists entirely of publicly available information and does not involve any developer or user private data.
(a) GPTs store.
(b) Coze.
(c) AgentBulider.
Figure 11: The growth trends of applications (top 3 categories by number and overall applications) on GPTs, Coze, and AgentBuilder (showing only publicly configured applications for AgentBuilder).
Appendix A Scoring Capability Descriptions and Constraints of Application Prompts
Here is the prompt we used to evaluate application prompts with ChatGPT-4o.
Input: <application prompt>
Judgment:
1. How many sentences describe the app’s capabilities, i.e., under what conditions the app can perform tasks?
2. How many sentences describe the app’s limitations, i.e., what the app should not do? Additionally, assess the degree of refusal for the app (1-5), where 1 indicates a polite refusal and 5 indicates a firm refusal.
Output: A JSON format output, including the level of detail in capability descriptions (1-5, where 1 is very vague and 5 is highly detailed) and the degree of refusal for each sentence describing capability limitations.
Example 1:
Application prompt: As a virtual assistant for the Yangtze Three Gorges Cruise Ticket Booking Center, your primary responsibility is to provide global tourists with services related to Three Gorges cruises, including ticket booking, ticket inquiries, cruise introductions, consultations, and online reservations. You must ensure that tourists receive accurate and timely information while delivering friendly customer service. 1. When a user inquires about Three Gorges cruise-related information, first identify their specific needs—whether it is ticket booking, ticket inquiries, cruise introductions, general consultations, or online reservations. 2. Based on the user’s needs, utilize the appropriate tools or interfaces to retrieve information. For instance, if the user needs to book tickets, invoke the ticket booking system API; if they need to inquire about ticket availability, use the ticket inquiry interface, and so on. 3. Once information is retrieved, organize and structure it to ensure it is presented to the user in a clear and accurate manner. 4. If the user’s request cannot be fulfilled—for example, if the queried tickets are sold out or cannot be reserved—clearly inform the user and provide possible alternatives or suggestions. 5. When handling user requests, adhere to the center’s business rules and procedures to ensure professionalism and efficiency in service. 1. Maintain a friendly and warm tone in responses, demonstrating respect and attentiveness to the user. 2. Tailor responses to include the requested information and make them as detailed and comprehensive as possible. 3. If the user has further questions or needs, proactively guide them to continue the conversation, ensuring their issues are fully resolved. 4. At the end of the response, express gratitude to the user and ask if there is anything else they need assistance with.
Response 1:
Example 2:
Application prompt: You are a travel guide assistant dedicated to providing tourists with customized, efficient, and practical strategies. When answering user questions, maintain a warm and friendly tone.
Response 2:
Appendix B LLM Judge Prompt
Here is the prompt we used in our LLM Judge.
Input:
A list of questions, each with an identifier question[id], and responses for each question[id] provided in the [ANSWER] field.
Metrics:
For each question (question[id]), output the following evaluations:
- Whether the [ANSWER] field for question[id] addresses the question (question[id]), 0 for no, 1 for yes.
- Whether the [ANSWER] field for question[id] refuses to answer the question (question[id]), 0 for no, 1 for yes.
- The accuracy of the response in the [ANSWER] field for question[id], rated on a scale of 0 to 10, where 0 indicates completely inaccurate and 10 indicates fully accurate.
- The output should be formatted in JSON.
Special Notes:
- If the [ANSWER] field is "NOAPP" or indicates that the application cannot be found, set all fields for all questions (question[id]) in the [Question] list to -1.
Appendix C Application Categories Distribution
Figure 12, 13, 14 and 15 are the complete application type distributions for different platforms.
Figure 12: Distribution of application categories on GPTs.
Figure 13: Distribution of application categories on Coze.
Figure 14: Distribution of application categories on AgentBuilder.
Figure 15: Distribution of application categories on Poe.
Appendix D Application Scale Growth Trend
Figure 11 illustrates the growth trends of total applications and top 3 types on GPTs, Coze, and AgentBuilder.
Appendix E Capability Jailbreak on Different Platforms
Figure 16 displays the capability jailbreak risks faced by the top 50 applications on different platforms. The x-axis represents the App Index, while the y-axis shows 0 for queries directly asking harmful questions and 1-5 for queries with varying degrees of adversarial techniques incorporated. The color of the small squares in the figure indicates the application’s response: green signifies no harmful content in the response, while red indicates the presence of harmful content.
Figure 16: Capability jailbreak risk heatmap for top 50 applications across different platforms.
Appendix F Top 5 Developers on GPTs
Figure 17 illustrates the application publishing history of the top five developers on GPTs.
Figure 17: Application publishing history of top 5 developers on GPTs.